# Realistic overground gait transitions are not sharp but involve gradually changing walk-run mixtures as per energy optimality

### v1

@@ Abstract
Humans use two qualitatively different gaits for locomotion, namely, walking and running — usually using walking at lower speeds and running at higher speeds. Researchers have examined when humans switch between walking and running on treadmills and have noted hystereses in these gait transition speeds. Here, we consider an ecologically realistic overground locomotion task, one requiring traveling a given long distance (800 meters or 2400 meters) in a prescribed time duration. Unlike on a treadmill, this task allows the human to change speed or gait during the trial to reach the destination on time: this task is akin to traveling to an appointment at a particular time from your office to another office, arriving neither early or late. We find that gait transition is not sharp, but instead involves a “gait transition regime” in which humans use a mixture of walking and running, using mostly walking at around 1.9 m/s and mostly running around 3.0 m/s — supporting earlier results over short distances (120 m). The presence of this gradually changing walk-run mixture is predicted by energy optimality. We hypothesize that this energy optimal behavior in this realistic overground conditions accounts for the hysteretic behavior in treadmill experiments, apparently switching earlier than predicted by energy optimality.
@@ 1 Introduction
Humans and many other terrestrial animals exhibit a number of different gaits [1, 2, 3, 4, 5]. Humans walk, run, and much more occasionally, skip [6, 7]. Horses walk, trot, canter, and gallop, and more occasionally, use other gaits [3, 8, 9]. Such gait transitions have most commonly been studied using treadmills (Figure 1). In these treadmill gait transition experiments, the treadmill speed is changed slowly, either in a continuous fashion with some fixed acceleration, or in a series of acceleration phases alternating with constant speed phases [10, 11, 12, 13, 14, 15, 16, 17, 18]. These experiments found that people switch between walking and running around 2 m/s, but sometimes, the walk to run transition speed was different from — and higher than — the run to walk transition speed [18, 11, 12] and the transition speeds were different from that predicted by energy optimality [10]. In such treadmill experiments, the gait transition is ‘sharp’, that is, happens at a particular speed where there is a preference of running over walking, or vice versa. Most humans do not spend their lives on treadmills, so their behavior may not already be energy optimal for such gait transition tasks without considerable learning [19, 20, 21]. Here, in contrast to these treadmill gait transition experiments, we show that overground gait transitions in realistic overground locomotion is more gradual and provide some clues for why there might exist distinct walk-to-run and run-to-walk speeds on a treadmill.
@@
Imagine you need to travel on foot from your home to an important appointment a kilometer away at a particular time (Figure 1). Unlike on a treadmill, where the speed is constrained, in this overground experiment, you can change speed or change gait. What might you do? If you start very early and have plenty of time, you might prefer to walk all the way. If you have very little time, you might need to run all the way. But if you had an intermediate amount of time, what might you do? Here, we perform this experiment for two long distances over 800 meters, and show that humans systematically use a mixture of walking and running when there is an intermediate amount of time. That is, we show that for such overground tasks, there is not a sharp gait transition speed below which walking is preferred and above which running is preferred. Having this mixture of walking and running instead of a sharp gait transition speed is energy optimal [1, 22, 23], and was earlier observed over short distance tasks in humans [1], so the primary experimental contribution of the current study is its demonstration over much longer distances.
@@
Figure 1: Experimental protocol. a) The classic treadmill-bound gait transition used in prior studies experiment does not provide any instantaneous choice in speed to the human, constraining speed at every moment, so that subjects are just asked to choose the gait they prefer at a given speed. b) Overground gait transition experiment used in the current study. Human subjects are asked to cover a given distance $D_{\mathrm{total}}$ in a fixed amount of time $T_{\mathrm{total}}$, as if they have to travel on foot to make an appointment, exactly arriving on time. Subjects carry a stop-watch counting down time remaining to prescribed duration and are free to choose and change their gait or speed whenever during each trial. Trials are repeated for different prescribed durations so as to achieve different average speeds.
@@ 2 Methods
Experiments were approved by the Ohio State Institutional Review Board and human subjects participated with informed consent. We had $N=14$ subjects (11 male, 3 female) with age 21.3 $\pm$ 2.9 years (mean $\pm$ s.d.), mass 70.35 $\pm$ 13.99 kg, height 1.80 $\pm$ 0.10 m, and leg length 0.91 $\pm$ 0.26 m. Subjects were instructed to travel one of two distances on foot, $D_{\mathrm{total}}=$ 800 m or 2400 m, in a pre-determined amount of time, $T_{\mathrm{total}}$, arriving exactly as allowed time expired, arriving neither early nor late (Figure 1b and Table 1). A stop-watch that counted down from $T_{\mathrm{total}}$ to zero was provided to the subject for reference throughout the trial. The $T_{\mathrm{total}}$ was not revealed to the subjects until just moments before starting a trial to eliminate as much prior planning as possible. The subjects received no additional instructions. The total distance and time constraints constrained only the average speed $v_{\mathrm{avg}}=D_{\mathrm{total}}/T_{\mathrm{total}}$, but not the instantaneous speeds throughout the trial, in contrast to a classic treadmill gait transition experiment (Figure 1a).
@@
For each total distance $D_{\mathrm{total}}$, we used four different $T_{\mathrm{total}}$ corresponding to four different average speeds: 1.92, 2.24, 2.62, and 2.98  ms-1. These speeds were chosen to be roughly in the treadmill gait transition regime [5, 24, 10, 11], as well as the regime in which humans used a mixture of walking and running in our earlier study where we examined this protocol for vastly shorter distances [1]. Because of the large distances involved here, each subject either performed two trials both at the 2400 m distance, or three trials, two at 800 m and one at 2400 m. In all, we performed $N_{\mathrm{trial}}=$ 42 total trials across the 14 subjects, with at least 9 trials per average speed. The trial assignment and trial order were randomized for each subject and across subjects. The total duration was not revealed to the subjects until a few moments before starting to eliminate as much pace planning as possible. All trials were performed on a 400 m outdoor track. Locomotion speeds were measured using GPS (10 Hz, VBOX Mini, Racelogic). Trials were video recorded and used to determine durations of walking and running during each trial, as visually classified by the investigators. The investigators classified the trial as running if there was a flight phase or if the hip moved downward and upward on a compliant-looking leg during stance phase, even if there did not seem to be a flight phase (jogging or grounded running); walking bouts are the complements of these running bouts [25, 26].
@@
Figure 2: Humans use a walk-run mixture. a) Fraction of time spent running for the 800 m trials (blue dots) and the 2400 m trials (orange dots). Fraction of running increases with average speed. b) Number of switches from running to walking or walking to running. The median running fraction (green line) and 25-75th percentile (light green band) are also shown in panels a-b. c) Histogram of time spent at various speeds for trials at each of the four average speeds. The histograms pool data over all subjects and all trials at the specific average speed. Each histogram has two peaks, one corresponding to walking (low speed) and another corresponding to running (high speed).
@@ 3 Results
Subjects typically completed the trials in exactly the prescribed amount of time, arriving neither too early nor too late, using walking and/or running, appropriately adjusting their speeds as needed. Only in 3 trials out of 42 trials did the subjects travelled so quickly initially that they had to be at virtually zero velocity for more than 10 sec, out of a typical trial duration of at least 540 seconds.
The subjects’ GPS-computed average speed was only 0.03 ms-1 different from the prescribed speed on average. The GPS-computed average speed and prescribed average speed were fit by the linear regression $v_{\mathrm{computed}}=0.99V_{\mathrm{avg}}+0.043$ ($p=10^{-37}$, $R^{2}=0.991$), which simultaneously demonstrates human ability to achieve different average speeds with high precision and the GPS-computed speed’s reliability (see acknowledgment regarding GPS accuracy).
@@
Subjects used a mixture of walking and running in 90% of the trials at the two intermediate speeds (2.22 and 2.6 ms-1). On average, walking dominates the walk-run mixture at the lower speeds and running dominates the walk-run mixture at the higher speeds (Figure 2a), so that the walk-run mixture gradually changes as speed is increased. The time fraction of walking decreases and the time fraction of running increases with average speed. Thus, for this overground gait transition task, there is no sharp gait transition speed but only a “gait transition regime”, which has substantial overlap with the speed range 2 m/s to 3 m/s. Despite the substantially greater distances involved (800 m and 2400 m), the overall trends in the walk-run fraction is similar to that observed for trials with much smaller distance (120 m) in the earlier study [1].
@@
In each of the walk-run mixtures, walking generally happens at low speeds and running generally happens at higher speeds. Figure 2c shows the time spent at various walking and running speeds at each of the four average speeds. Each histogram shows two peaks, one at a walking speed and another at a running speed. Considering these histograms (Figure 2c) with the time fraction of walking versus running (Figure 2a) suggests that people accomplish the required average speed by averaging over a slow walking speed and a faster running speed, as qualitatively predicted by energy optimality [1]. As the average speed increases, the speed distribution changes so that there is more time spent at the higher running speeds, resulting in a larger peak at the higher speed. Speed 1.98 ms-1 is dominated by walking and speed 2.96 ms-1 is dominated by running.
@@
Humans typically switch gait only a small number of times even for long distances lowering switching costs.
@@
Pure energy optimality in the absence of fatigue or any uncertainty predicts that in the regime where a walk-run mixture is optimal, there should be exactly one switch between walking and running. Multiple switches between walking and running is not optimal. Here, we find that the median number of switches is one for three out of the four speeds we considered, with speed 2.62 m/s having two switches as median, with some variability around this median (Figure 2b). If we assume the cost of a single switch to be $m\left(V_{\mathrm{run}}^{2}-V_{\mathrm{walk}}^{2}\right)/2$ approximately [27, 28], we estimate about 3.4 J per unit body mass when the walking speed is 1.5 m/s and the running speed is 3 m/s. This gait and speed switch cost of 3.4 J/kg is negligible compared to about 2300 J/kg for walking 800 m at 1.5 m/s, using steady walking costs from [29, 30]. Despite this cost per switch being negligible, humans only switch gaits a small number of times, still keeping the switching cost negligible.

### v2

# Overground gait transitions are not sharp but involve gradually changing walk-run mixtures even over long distances

@@ Abstract
Humans typically walk at low speeds and run at higher speeds. Previous studies of transitions between walking and running were mostly on treadmills, but real-world locomotion allows more flexibility. Here, we study overground locomotion over long distances (800 m or 2400 m) under time constraints, simulating everyday scenarios like going to an appointment. Unlike on treadmills, participants can vary both speed and gait during this task. We find that gait transition in this overground task occurs over a broad ‘gait transition regime’ spanning average speeds from 1.9 m/s to 3.0 m/s. In this regime, people use mixtures of walking and running: mostly walking at low average speeds (around 1.9 m/s) and mostly running at high average speeds (3.0 m/s); the walk vs run fraction gradually changes between these speed limits. Within any walk-run mixture, there is a speed gap between the walking and running. These gait mixtures and their specific structure are predicted by energy optimality. These findings extend earlier results from much shorter distance tasks, showing that similar energetic principles govern longer, more physically and cognitively demanding tasks. Overall, our results highlight the role of whole-task energy minimization including transients in shaping human locomotion.
@@ 1 Introduction
Humans and many other terrestrial animals exhibit a number of different gaits [1, 2, 3, 4, 5]. Humans walk, run, and much more occasionally, skip [6, 7]. Horses walk, trot, canter, and gallop, and more occasionally, use other gaits [3, 8, 9]. Transitions between such gaits have most commonly been studied using treadmills. In these treadmill gait transition experiments (Figure 1), the treadmill speed is changed slowly, either in a continuous fashion with some fixed acceleration, or in a series of acceleration phases alternating with constant speed phases [10, 11, 12, 13, 14, 15, 16, 17, 18]. These experiments found that people switch between walking and running around 2 m/s, but sometimes, the walk to run transition speed was different from — and higher than — the run to walk transition speed [18, 11, 12]. Importantly, the transition speeds were different from that predicted by energy optimality [10, 19], even though there is extensive evidence indicating energetics as a determinant of locomotor choice in numerous other settings [20, 21, 22, 23, 24, 25, 26, 2, 27, 23, 28, 29]. In such treadmill experiments, the gait transition is ‘sharp’, that is, happens at a particular speed where there is a preference of running over walking, or vice versa. Of course, humans do not spend their lives on treadmills, so their behavior may not already be energy optimal for such gait transition tasks without considerable learning and experience near the gait transition speeds [30, 31, 32]. Here, in contrast to these treadmill gait transition experiments, we show that overground gait transitions in realistic overground locomotion is more gradual and provide some clues for why there might exist distinct walk-to-run and run-to-walk speeds on a treadmill.
@@
Say you need to travel on foot from your home to an important appointment a kilometer away at a particular time, which provides a time-deadline (Figure 1). Unlike on a treadmill, where the speed is constrained, in this overground experiment, you can change speed or change gait. If you start very early and have plenty of time, you might prefer to walk all the way. If you have very little time, you might need to run all the way. But if you had an intermediate amount of time, what might you do? Here, we perform this experiment for two long distances, 800 m and 2400 m, and show that humans systematically use a mixture of walking and running when there is an intermediate amount of time. That is, we show that for such overground tasks, there is not a sharp gait transition speed below which walking is preferred and above which running is preferred. Having this mixture of walking and running instead of a sharp gait transition speed is energy optimal [1], and was earlier observed in humans by Long and Srinivasan [1] in a much shorter distance version of this task. So, the current manuscript can be considered an expansion of this previous study [1], with the experimental contribution of the current study being consideration of much longer distances, but with additional contributions in terms of data analyses and computational work not present in the previous study [1]. Longer distances make the energy costs more salient to the human, while also making the task of adjusting speed and gait in an optimal manner more cognitively demanding. So, it is a priori unclear whether humans will use walk-run mixtures as they did over short distances. This new data over longer distances also allows us to characterize how the preferred walking and running speeds, the fractions of walking and running, and the within-trial gait transition speed vary as a function of the mean locomotion speed, not previously characterized. We then show that these trends in the human walk-run mixtures are also qualitatively consistent with energy minimization.
@@
Figure 1: Experimental protocol. a) The classic treadmill-bound gait transition used in prior studies experiment does not provide any instantaneous choice in speed to the human, constraining speed at every moment, so that participants are just asked to choose the gait they prefer at a given speed. b) Overground gait transition experiment used in the current study. Human participants are asked to cover a given distance $D_{\mathrm{total}}$ in a fixed amount of time $T_{\mathrm{total}}$, as if they have to travel on foot to make an appointment, exactly arriving on time. Participants carry a stop-watch counting down time remaining to prescribed duration and are free to choose and change their gait or speed whenever during each trial. Trials are repeated for different prescribed durations so as to achieve different average speeds.
@@ 2 Methods
Experiments were approved by the Ohio State Institutional Review Board and human participants took part in them with informed consent. We had $N=14$ participants (11 male, 3 female) with age 21.3 $\pm$ 2.9 years (mean $\pm$ s.d.), mass 70.35 $\pm$ 13.99 kg, height 1.80 $\pm$ 0.10 m, and leg length 0.91 $\pm$ 0.26 m. participants were instructed to travel one of two distances on foot, $D_{\mathrm{total}}=$ 800 m or 2400 m, in a pre-determined amount of time, $T_{\mathrm{total}}$, arriving exactly as allowed time expired, arriving neither early nor late (Figure 1b and Table 1). A stop-watch that counted down from $T_{\mathrm{total}}$ to zero was provided to the participant for reference throughout the trial. The $T_{\mathrm{total}}$ was not revealed to the participants until just moments before starting a trial to eliminate as much prior planning as possible. The participants received no additional instructions. The total distance and time constraints constrained only the average speed $v_{\mathrm{avg}}=D_{\mathrm{total}}/T_{\mathrm{total}}$, but not the instantaneous speeds throughout the trial, in contrast to a classic treadmill gait transition experiment (Figure 1a).
@@
For each total distance $D_{\mathrm{total}}$, we used four different $T_{\mathrm{total}}$ corresponding to four different average speeds: 1.92, 2.24, 2.62, and 2.98  ms-1. These speeds were chosen to be roughly in the treadmill gait transition regime [5, 33, 10, 11], as well as the regime in which humans used a mixture of walking and running in our earlier study where we examined this protocol for a shorter distance [1]. Because of the large distances involved here, each participant either performed two trials both at the 2400 m distance, or three trials, two at 800 m and one at 2400 m. In all, we performed $N_{\mathrm{trial}}=$ 42 total trials across the 14 participants, with at least 9 trials per average speed. The trial assignment and trial order were randomized for each participant and across participants. The total duration was not revealed to the participants until a few moments before starting to eliminate as much pace planning as possible. All trials were performed on a 400 m outdoor track. Locomotion speeds were measured using GPS (10 Hz, VBOX Mini, Racelogic) and was available for all but one subject for whom this data was corrupted. Trials were video recorded and used to determine durations of walking and running during each trial, as visually classified by the investigators. The investigators classified the trial as running if there was a flight phase or if the hip moved downward and upward on a compliant-looking leg during stance phase, even if there did not seem to be a flight phase (jogging or grounded running); walking bouts are the complements of these running bouts [34, 35].
Computational model.
Our energy-based account of these overground gait transitions is based on minimizing the energy cost of the entire locomotor task, including the transients like acceleration, deceleration, and gait transitions — as opposed to minimizing something instantaneous or steady-state like the cost of transport [36, 5, 22]. Earlier work had shown that this whole-task energy optimality framing explains locomotor phenomena better [1, 21, 20], with some recent work finding additional evidence [37].
For each trial, given travel distance $D_{\mathrm{total}}$ and time duration $T_{\mathrm{total}}$, we compute the energy optimal walking speed $V_{\mathrm{walk}}$ and running speeds $V_{\mathrm{walk}}$, and the respecting walking and running durations ($T_{\mathrm{walk}}$ and $T_{\mathrm{run}}$). The individual gait durations must satisfy the total duration constraint $T_{\mathrm{walk}}+T_{\mathrm{run}}=T_{\mathrm{total}}$, and along with the speeds, they satisfy the total distance constraint $V_{\mathrm{walk}}T_{\mathrm{walk}}+V_{\mathrm{walk}}T_{\mathrm{walk}}=D_{\mathrm{total}}$. The total energy cost of a bout is computed as the sum of the walking energy cost, running bout energy cost, and the cost for the transitions between gaits as well as the changes in speed [21, 38, 1]. The metabolic rate for walking and running as a function of speed are of the form $\dot{E}=a_{0}+a_{1}V+a_{2}V^{2}$, with different coefficients [1, 39, 40], the cost for changing speed is proportional to the kinetic energy change [21], and the gait change cost is equal to that in [38].
These calculations are identical to that in [1] specialized to the task conditions herein, and are performed using a nonlinear optimization procedure. Here, in addition, noting that the energy landscape tends to have relatively flat minima [21, 41, 20, 31], we provide a sensitivity analysis: we compute the set of all speeds and gait durations that are within 1% and 2% of the previously computed optimal energy cost. To do this, we scan over a range of walking time durations for each speed, and solve for the optimal walking and running speeds for each such walking duration along with the corresponding energy costs, which then allows us to characterize the speeds and durations within 1% or 2% from the optimal energy cost.
@@
Figure 2: Humans use a walk-run mixture. a) Fraction of time spent running for the 800 m trials (blue dots) and the 2400 m trials (orange dots). Fraction of running increases with average speed. b) Number of switches from running to walking or walking to running. The median running fraction (green line) and 25-75th percentile (light green band) are also shown in panels a-b. c) Histogram of time spent at various speeds for trials at each of the four average speeds. The histograms pool data over all participants and all trials at the specific average speed. Each histogram has two peaks, one corresponding to walking (low speed) and another corresponding to running (high speed).
@@ 3 Results
Participants typically completed the trials in exactly the prescribed amount of time, arriving neither too early nor too late, using walking and/or running, appropriately adjusting their speeds as needed. Only in 3 trials out of 42 trials did the participants travelled so quickly initially that they had to be at virtually zero velocity for more than 10 sec, out of a typical trial duration of at least 540 seconds.
The participants’ GPS-computed average speed was only 0.03 ms-1 different from the prescribed speed on average. The GPS-computed average speed and prescribed average speed were fit by the linear regression $v_{\mathrm{computed}}=0.99V_{\mathrm{avg}}+0.043$ ($p=10^{-37}$, $R^{2}=0.991$), which demonstrates human ability to achieve different average speeds with high precision.
Figure 3: a) Distribution of walking and running speeds changes gradually with average speed, with more time spent walking at low speeds and more time spent running at high speeds. Running speeds are systematically higher than walking speeds at each average speed. The overlap observed between walking and running speeds is largely an artifact of pooling participants (see Supplementary Figure S1). b) Gait transition speeds obtained from a logistic classifier do not change significantly speed. c) Mean running speeds increase with average speed. d) Mean walking speed remains largely constant or slowly decreases.
Figure 4: a) Optimal walking and running fractions as a function of average speed, assuming the optimization has perfect knowledge and has zero error. During the walk-run mixture regime, the running speed slowly increases and the walking speed remains roughly constant. b) Optimal walk-run fractions and optimal speeds, allowing that the metabolic energy can be within 1% (darker band) or 2% (lighter band) of the optimal metabolic cost.
@@
Participants used a mixture of walking and running in 90% of the trials at the two intermediate speeds (2.22 and 2.6 ms-1). On average, walking dominates the walk-run mixture at the lower speeds and running dominates the walk-run mixture at the higher speeds (Figure 2a), so that the walk-run mixture gradually changes as speed is increased. The time fraction of walking decreases and the time fraction of running increases with average speed. Thus, for this overground gait transition task, there is no sharp gait transition speed but only a “gait transition regime”, which has substantial overlap with the speed range 2 m/s to 3 m/s. Despite the substantially greater distances involved in this study (800 m and 2400 m), the overall trends in the walk-run fraction is similar to that observed for trials with much smaller distance (120 m) in the earlier study [1], suggesting that this walk-run mixture is a natural behavior for this task.
@@
The speed distributions during each trial have two peaks (Figure 2b). After categorizing every moment as either walking or running, we obtain separate walking speed and running speed histograms for each trial. These gait-specific histograms illustrate that in each of the walk-run mixtures, walking generally happens at low speeds and running generally happens at higher speeds (Figure 3a; see Supplementary Figure S1 for individual trial histograms). Specifically, people accomplish the required average speed by averaging over a slow walking speed and a faster running speed, as qualitatively predicted by energy optimality [1]. As the average speed increases, the speed distribution changes so that there is more time spent at the higher running speeds, resulting in a larger peak at the higher speed: speed 1.98 ms-1 is dominated by walking and speed 2.96 ms-1 is dominated by running.
For most participants that used a mixture of walking and running, there is typically no overlap in the speeds used respectively for walking and running: typically there is a substantial gap in the speeds used. This gap is not obvious in the gaitwise histograms pooled across subjects (figure 3a) but is clear from visual inspection of the individual subjects’ gaitwise speed histograms (appendix figure S1). Thus, subjects do not use intermediate speeds, instead jumping directly from walking speeds to higher running speeds.
Systematic changes in running speed, but not walking speed or gait transition speed.
In this gait transition regime, mean running speed increased systematically with average speed, thereby contributing to the increased average speed (figure 3b-c). In contrast, mean walking speeds remained roughly constant as average speed increased. We obtained a ‘within-trial gait transition speed’ for each trial by applying a logistic classifier on the walking and running speed data and found that this gait transition speed also does not change systematically with speed, instead remaining roughly constant (figure 3b-c).
Predictions from energy optimality agree qualitatively with experiments.
Computational predictions from energy optimality show that it is optimal to use walking alone for locomotion tasks with low average speeds (too much time), it is optimal to use running alone for tasks with high average speeds (very little time), and it is optimal to use a walk-run mixture for intermediate optimal speeds, as observed here (figure 4a,b). In the walk-run mixture regime, the model predicts that the walking fraction is high for low average speeds and the running fraction is high for high average speeds. Further, the model predicts that in the walk-run mixture regime, the walking speeds remains nearly constant but the running speed increases with average speed (figure 4a,b), as qualitatively observed in experiments (figure 3b-c). The running speed increasing with average speed requires there to be a cost for speed transitions [21], as not having any transition costs results in constant walking and running speeds across the different average speeds (see [1]). This result reemphasizes the importance of such transient costs in predicting human behavior.
Our sensitivity analysis reveals that the set of speeds and durations within 1% or 2% of the optimal energy costs is relatively broad, revealing a flat energy landscape, allowing humans high variability in speed with little energy penalty (figure 4b-c). This analysis shows that the energy cost is less sensitive to the running speed (larger band of speeds within 1% or 2% of optimality) compared to the walking speed; this lower sensitivity is inherited from the lower curvature of the running metabolic rate [5], so even normal running cost of transport is only weakly sensitive to running speed [39].
@@
Humans typically switch gait only a small number of times even for long distances.
@@
Pure energy optimality in the absence of fatigue or any uncertainty predicts that in the walk-run mixture regime, there should be exactly one switch between walking and running; multiple switches between walking and running would require more speed and gait transition costs. For three out of the four speeds considered here, the median number of switches is one; speed 2.62 m/s has two switches as median, with some variability around this median (Figure 2b). If we assume the cost of a single switch to be $m\left(V_{\mathrm{run}}^{2}-V_{\mathrm{walk}}^{2}\right)/2$ approximately [21, 38], we estimate about 3.4 J per unit body mass when the walking speed is 1.5 m/s and the running speed is 3 m/s. This gait and speed switch cost of 3.4 J/kg is negligible compared to about 2300 J/kg for walking 800 m at 1.5 m/s, using steady walking costs from [40, 22]. Despite this cost per switch being negligible, humans only switch gaits a small number of times, still keeping the switching cost negligible.